Assuming 100% efficient energy conversion, how much water stored behind a 50 centimeters high hydroelectric dam would be required to charge the battery?
1
Expert's answer
2020-06-10T18:24:59-0400

The power of the dam is


"P = \u03c1ghQ,"

where "Q" - flow of water in m3/s.

Assume that the battery has "V=12 \text{ V}, q=50\text{ A}\cdot\text{h}." Its power will be


"P=\frac{qV}{t}."

Equate:


"ρghQ=\frac{qV}{3600\text{ s}},"

the flow of water will be


"Q=\frac{qV}{3600\rho gh}=\frac{50\cdot 12}{3600\cdot1000\cdot9.8\cdot0.5}=\frac{0.12\text{ m}^3}{3600\text{ s}},"

the total volume for 1 hour is


"v=Q\cdot3600=0.12 \text{ m}^3."